DIODE LASER SPECTROSCOPY OF SD_{3}^{+}

\underline{MARIA M. SANZ}, CHANGHONG XIA, and STEPHEN C. FOSTER, Department of Chemistry, Mississippi State University, Mississippi State, MS 39762-9573.

Three of the four fundamental bands of the SH_{3}^{+} ion were observed by Amano et al.^{a} Much less is known about the fully-deuterated isomer, SD_{3}^{+}. An ab initio calculation by Botschwina et al.^{b} predicted the ν_{1} and ν_{3} band origins at 1832 and 1839 cm^{−1} respectively, but there have been no previous experimental observations of this isomer.

In this talk we will discuss our observation of the absorption spectrum of SD_{3}^{+} in the region between 1750-1870 cm^{−1}. The ion was generated in a methanol-cooled hollow cathode discharge of D_{2}, containing a trace quantity of OCS, and a series of spectra were recorded by means of a diode laser system. The ν_{1} and ν_{3} fundamentals of this ion perturb each other via a Coriolis interaction and a full analysis requires us to simultaneously fit both bands. The analysis is currently in progress, and our results will be discussed at this meeting.

aT. Nakanaga and T. Amano, Chem. Phys. Lett. 134 (2) , 195 (1987). T. Amano, K. Kawaguchi and E. Hirota, J. Mol.

Spectrosc. 126, 177 (1987). T. Nakanaga and T. Amano, J. Mol. Spectrosc. 133, 201 (1989)

bP. Botschwina, A. Zilch, H.-J. Werner, P. Rosmus and E.-A. Reinsch, J. Chem.Phys. 85 (9), 5107 (1986)